OpenStudy (anonymous):

Solve (1/216)^3nx1/36=1/6

OpenStudy (anonymous):

\[x=\frac{ 60466176}{n}\]

OpenStudy (anonymous):

Can you give me step by step?

OpenStudy (anonymous):

(n x)/362797056 = 1/6 Divide both sides by n/362797056....................you get x = 60466176/n

OpenStudy (anonymous):

Im really confused.

OpenStudy (anonymous):

k wait draw your equation exactly

OpenStudy (anonymous):